# Trans-dimensional Hamiltonian model selection and parameter estimation from sparse, noisy data

### v1

@@ 5.2.1 Performance in sparse data limit
To evaluate the performance of our recovery method in the limit of sparse data, we varied the number of $\tau_{j}$ sampled from the coherence signals from a simulated spin bath of ten nuclear spins (Fig. 3a). We then assessed whether the posterior distributions of the spin configurations (Fig. 3b) and relative decoherence (Fig. 3c) obtained from the hybrid algorithm described in Sec. 3.4 successfully captured the simulated nuclear spin configurations and relative decoherence parameter $\lambda$. To ensure these results generalize across spin configurations, we analyzed the recovery rate for the hyperfine couplings, the detection rate for the number of spins, and the false positive rate for nuclear detection for 16 different simulated spin baths containing 5 to 20 spins, with a hyperfine magnitude greater than 5 kHz (Fig. 3d).
For the sample bath consisting of 10 nuclear spins and varying number of $\tau_{j}$, we observe, as expected, that the fidelity of the recovered signal improves with an increased number of sampled $\tau_{j}$ (Fig. 3a). However, what we are interested in is not necessarily to best recover the underlying signal but to recover the underlying nuclear spin bath configuration that gives rise to the measured signal. We find that even for signals sampled as sparsely as 50 $\tau_{j}$, the posterior distribution of hyperfine couplings of spins contains all ten of the simulated spins (Fig. 3b), and the signal sampled as sparsely as 25 $\tau_{j}$ recovers the nuclear spin with the strongest hyperfine magnitude. As expected, the more sparsely sampled signals result in more fictitious spins in the posterior compared to more densely sampled spins. The density of the time-point sampling is related to the ill-posedness of the inverse problem; from this sample bath, we also observe that the ability to resolve hyperfine couplings from sparse signals is highly correlated with the strength of the hyperfine magnitude. That is, the strongly coupled spins are clearly resolved in the posterior distributions in Fig. 3b, whereas we are unable to clearly resolve the more weakly coupled spins are unable, even at relatively dense samplings of the signals. We emphasize that the failure to resolve weakly coupled spins is related to the inherence ill-posedness of the inverse problems. Weakly coupled spins lead to smaller dips in the fringes of the coherence signal; as a result, many different combinations of weakly coupled spins can give rise to identical patterns of peaks in the coherence signal, making it impossible to resolve individual weakly coupled spins from a single sparse coherence signal originating from a dynamical decoupling experiment.
@@
The accuracy of recovery rates is notably lower for weakly coupled nuclei ($<25$ kHz) across multiple metrics, including detection rate, number of recovered spins, and false positive rate, as seen in Fig. 3d-f. This behavior arises from the inherent ill-posedness of the problem in the sparse data regime: distinct configurations of weakly coupled spins can generate identical coherence signals with comparable likelihoods, limiting the ability to resolve individual weakly-coupled nuclear spins. This limitation is a fundamental consequence of the ill-posedness of this inverse problem in the sparse data regime rather than a failure of the hybrid algorithm to accurately resolve the simulated nuclear spin configuration.
@@ 5.3 Validation on experimental data
Figure 5: Validation of hybrid MCMC approach on experimental data: (a) Average error per data point as a function of the number of MCMC steps, with five random ensembles initialized for 25,000 steps each; the first 10,000 steps are treated as burn-in, and the remaining 15,000 steps are the posterior spin configurations. (b) Experimental coherence data (250 data points from 6 $\mu$s to 8 $\mu$s of a 32-pulse CPMG sequence) to we applied the hybrid MCMC algorithm, and the reconstructed signal using the best-fit spin configuration from the posterior distribution. (c) Posterior distribution of the total number of spins in the recovered spin configurations. (d) Posterior distribution of hyperfine couplings, with the 50 reported nuclear spin hyperfine couplings from the reference study overlaid. (e) Posterior distribution of the number of spins by spin magnitude ($\sqrt{A_{\perp}^{2}+A_{\parallel}^{2}}$), with the mode displayed and the reference value reported in parentheses.
We validate the hybrid MCMC approach by applying it to coherence data collected from a single NV center at a temperature of 3.7 K using a 32-pulse CPMG pulse sequence, with $\tau$ ranging from 6 to 8 $\mu$s. For more details on experimental procedure, see App. C. This NV center has well-characterized hyperfine interactions and lattice positions of surrounding nuclear spins, determined from high-resolution multidimensional spectroscopy and correlated sensing experiments [16, 35]. These experimentally determined values serve as reference values.
We use the hyperfine couplings and lattice positions calculated in [34] as theoretical reference values. The alignment procedure between these theoretical values and the experimentally determined values is described in Appendix F. The hybrid MCMC algorithm (Alg. 4) was initialized with five independent ensembles, each comprising 25,000 MCMC steps. All ensembles converged to the same error regime within 5,000 steps (Fig. 5a), and the initial 10,000 steps were discarded as burn-in. The posterior distribution was constructed from the remaining 15,000 steps.
@@
This hybrid formulation enables the inference algorithm to capture both local pointwise agreement and global distributional similarity. The squared-error term ensures that the model fits the observed decay values at each delay time $\tau_{i}$, while the Wasserstein term penalizes global mismatches in the overall structure of the coherence decay curve—an important feature when dealing with noisy, underdetermined, or aliased experimental data [36]. This additional regularization is not required when testing our hybrid MCMC approach on simulated data generated from known hyperfine parameters. In the simulation case, the model is sampled from the same distribution as the simulated data, and there is no error due to mismatches between the model and physical reality. Therefore, a standard likelihood based on squared residuals is sufficient for an accurate reconstruction, in the case of the simulated data.
In contrast, when analyzing experimental data, we know that the true hyperfine couplings differ systematically from the theoretical values used to define the sampling space [34]. As a result, directly comparing the predicted and experimental signals point-wise can lead to overfitting or biased inference. Incorporating the Wasserstein distance allows the inference process to remain robust to such mismatches by focusing on the agreement in the shape and distribution of the coherence signal, rather than on the exact point-wise correspondence. The weight $\zeta=0.5$ was selected after benchmarking several values, as detailed in Appendix E, where we show that this value of $\zeta$ balances local fit and global structure matching.
@@
The posterior distribution of the total number of nuclear spins (Fig. 5c), with a modal value of 46, shows excellent agreement with the experimental reference value of 50. We further evaluated the posterior distribution of hyperfine couplings (Fig. 5d). A total of 45 out of the 48 experimentally identified nuclear spins are detected in the posterior distribution, with 22 of these spins exhibiting detection rates exceeding their baseline probabilities (Table LABEL:tab:hf_compare). The calculation of detection rates and baseline probabilities is detailed in Appendix F.
@@
In contrast, our method achieves comparable performance using only 250 data points from a single 32-pulse experiment and runs in approximately 8 hours using only 5-cores of a single CPU node (Intel Xeon E5-2680 v4), without requiring pre-training or additional tuning when experimental conditions are changed. Compared to another non-ML approach proposed in Ref. [37], which identified 14 experimental spins using high-resolution data (around 10,000 time points of the coherence signal) across multiple 13C Larmor periods, our method identifies more spins using a much sparser dataset. The tradeoff of working with sparse data is the persistence of ill-posedness; however, our Bayesian inference framework reduces the number of plausible lattice sites from 3,518 to 230 by leveraging posterior detection probabilities.
@@ 6 Discussion
A potential limitation of the hybrid MCMC approach arises when the true Hamiltonian does not lie within the predefined class of candidate models. However, the framework is naturally compatible with established techniques for model selection and averaging, which can mitigate issues related to model misspecification. Bayesian model averaging [48], mixture models [49], and ensemble strategies [50] such as bagging offer systematic approaches to incorporate uncertainty over model structure. In particular, bagging methods—originally developed in the context of machine learning—construct ensembles by training models on bootstrapped subsets of the data and combining their predictions through voting, averaging, or regularized aggregation to improve robustness; such procedures are adopted in sparse data regimes where small variations in the dataset may lead to significant shifts in the inferred posterior [51]. By enabling principled treatment of both parameter and model uncertainty, such extensions could preserve the interpretability and physical grounding of the Bayesian approach while enhancing its stability and generalizability.
@@ Acknowledgments
We thank F. Joseph Heremans, Mykyta Onizhuk, Daniel Sanz-Alonso, Nathan Waniorek, Christopher Egerstrom, and Benjamin Pingault for useful discussions. A.N.P. acknowledges support from the DOE CSGF under Award Number DE-SC0022158. This research used resources of the University of Chicago Research Computing Center.
@@ Appendix A Hyperparameter dependence
Figure 6: Error trajectories of 5 ensembles as a function of walker steps for varying $\sigma^{2}$. The recovery algorithm was applied to the same coherence signal generated using $N=16$ CP pulses, $\varepsilon=0.001$ shot noise, hyperfine perturbation $0.001$, $\tau_{\text{max}}=0.008$ ms, and 250 interpulse times sampled from 15 simulated spins with $R_{\text{spin}}=5$.
@@
We find that the hybrid algorithm presented is highly robust to $R_{\text{spin}}$, but the performance of the hybrid algorithm is sensitive to $\sigma^{2}$. In Fig. 7, we find that the variance, mean, and convergence properties of the error trajectories are uncorrelated with the size of the walker radius over the lattice sites $R_{\text{spin}}$. That is, the probabilistic selection rules introduced by the hybrid algorithm bias the walkers toward the same solution, regardless of the size of the neighborhood by which each lattice walker is constrained. Notably, there is a physical limit on how small the walker radius can be, which is the distance of nearest neighbors between lattice sites. For diamond, the distance between nearest neighbors is around $1.54$. Unless otherwise stated, all simulations and experimental data recovery were performed with $R_{\text{spin}}=5$.
@@ Appendix B Performance of hybrid algorithm in dense, low noise limit
Only other previous non-ML method is algorithmic decomposition, relies on fitting a line to points generated by the peaks as a function of the number of Larmor periods, we are interested here in data that may only contain two or three Larmor periods, and it is not good form to fit a line to two or three points.

### v2

@@ 5.2.1 Performance in sparse data limit
To evaluate the performance of our recovery method in the limit of sparse data, we vary the number of $\tau_{j}$ sampled from the coherence signals from a simulated spin bath of ten nuclear spins (Fig. 3a). We then assess whether the posterior distributions of the spin configurations (Fig. 3b) and relative decoherence (Fig. 3c) obtained from the hybrid algorithm described in Sec. 3.4 successfully capture the simulated nuclear spin configurations and relative decoherence parameter $\lambda$. To ensure these results generalize across spin configurations, we analyze the recovery rate for the hyperfine couplings, the detection rate for the number of spins, and the false positive rate for nuclear detection for 16 different simulated spin baths containing 5 to 20 spins, with a hyperfine magnitude greater than 5 kHz (Fig. 3d).
For the sample bath consisting of 10 nuclear spins and varying number of $\tau_{j}$, we observe, as expected, that the fidelity of the recovered signal improves with an increased number of sampled $\tau_{j}$ (Fig. 3a). However, we are not necessarily interested in reconstructing the coherence signal, but rather we seek to recover the underlying nuclear spin bath configuration that gives rise to the observed coherence signal. We find that even for signals sampled as sparsely as 50 $\tau_{j}$, the posterior distribution of hyperfine couplings of spins contains all ten of the simulated spins (Fig. 3b), and the signal sampled as sparsely as 25 $\tau_{j}$ recovers the nuclear spins with the strongest hyperfine magnitudes. As expected, the more sparsely sampled signals result in more fictitious spins in the posterior compared to more densely sampled spins. The density of the time-point sampling is related to the ill-posedness of the inverse problem; from this sample bath, we also observe that the ability to resolve hyperfine couplings from sparse signals is highly correlated with the strength of the hyperfine magnitude. That is, the strongly coupled spins are clearly resolved in the posterior distributions in Fig. 3b, whereas we are unable to clearly resolve the more weakly coupled spins, even at relatively dense samplings of the signals. We emphasize that the failure to resolve weakly coupled spins is related to the inherence ill-posedness of the inverse problems. Weakly coupled spins lead to smaller dips in the fringes of the coherence signal; as a result, many different combinations of weakly coupled spins can give rise to identical patterns of peaks in the coherence signal, making it impossible to resolve individual weakly coupled spins from a single sparse coherence signal originating from a dynamical decoupling experiment.
@@
The accuracy of recovery rates is notably lower for weakly coupled nuclei (<25<25 kHz) across multiple metrics, including detection rate, number of recovered spins, and false positive rate, as seen in Fig. 3d-f. This behavior arises from the inherent ill-posedness of the problem in the sparse data regime: distinct configurations of weakly coupled spins can generate identical coherence signals with comparable likelihoods, limiting the ability to resolve individual weakly-coupled nuclear spins. This limitation is a fundamental consequence of the ill-posedness of this inverse problem in the sparse data regime rather than a failure of the hybrid algorithm to accurately resolve the simulated nuclear spin configuration.
@@ 5.3 Validation on experimental data
[…]
We use the hyperfine couplings and lattice positions calculated in [34] as theoretical reference values. The alignment procedure between these theoretical values and the experimentally determined values is described in Appendix F. The hybrid MCMC algorithm (Alg. 4) is initialized with five independent ensembles, each comprising 25,000 MCMC steps. All ensembles converged to the same error regime within 5,000 steps (Fig. 5a), and the initial 10,000 steps were discarded as burn-in. The posterior distribution was constructed from the remaining 15,000 steps.
@@
This hybrid formulation enables the inference algorithm to capture both local pointwise agreement and global distributional similarity. The squared-error term ensures that the model fits the observed decay values at each delay time $\tau_{i}$, while the Wasserstein term penalizes global mismatches in the overall structure of the coherence decay curve—an important feature when dealing with noisy, underdetermined, or aliased experimental data [36]. This additional regularization is not required when testing our hybrid MCMC approach on simulated data generated from known hyperfine parameters. In the simulation case, the model is sampled from the same distribution as the simulated data, and there is no error due to mismatches between the model and physical reality. Therefore, a standard likelihood based on squared residuals is sufficient for an accurate reconstruction, in the case of the simulated data. In contrast, when analyzing experimental data, we know that the true hyperfine couplings differ systematically from the theoretical values used to define the sampling space [34]. As a result, directly comparing the predicted and experimental signals point-wise can lead to overfitting or biased inference. Incorporating the Wasserstein distance allows the inference process to remain robust to such mismatches by focusing on the agreement in the shape and distribution of the coherence signal, rather than on the exact point-wise correspondence. The weight $\zeta=0.5$ was selected after benchmarking several values, as detailed in Appendix E, where we show that this value of $\zeta$ balances local fit and global structure matching.
@@
The posterior distribution of the total number of nuclear spins (Fig. 5c), with a modal value of 46, shows excellent agreement with the experimental reference value of 50. We further evaluated the posterior distribution of hyperfine couplings (Fig. 5d). A total of 45 out of the 48 experimentally identified nuclear spins are detected in the posterior distribution, with 22 of these spins exhibiting detection rates exceeding their baseline probabilities (Table LABEL:tab:hf_compare). The calculation of detection rates and baseline probabilities is detailed in Appendix F. One strongly coupled spin (labeled C36 in Table 2 with ($A_{\parallel},A_{\perp}$) = (84.52, 114.42) kHz) is not detected in the posterior distribution (see Fig. 5d). This illustrates the combined effects of sparse experimental data and known discrepancies between DFT-computed and experimentally measured hyperfine couplings: when the theoretical hyperfine values deviate from experiment, the posterior detection probability can be reduced even for otherwise well-resolved spins. Appendix D provides a detailed analysis showing that recovery rates for both spin number and hyperfine couplings remain high when the input couplings are accurate to within $\sim 1$ kHz, while larger discrepancies between the DFT-computed and experimentally-measured hyperfine couplings, as in the case of C36, lead to decreased detection.
@@
In contrast, our method achieves comparable performance using only 250 data points from a single 32-pulse experiment and runs in approximately 8 hours using only 5-cores of a single CPU node (Intel Xeon E5-2680 v4), without requiring pre-training or additional tuning when experimental conditions are changed. The experimental dataset used here was originally acquired for different studies [16, 35]; we sub-sampled the available time points to test the inference framework in a regime currently inaccessible to prior inverse-problem techniques, demonstrating that accurate spin detection can be achieved with substantially fewer measurements. Typical acquisition of a single 32-pulse coherence signal with 250 time points requires roughly 5–10 hours under standard laboratory conditions, including NV selection, alignment, and calibration, making the measurement time comparable to the runtime of the inference algorithm. Compared to another non-ML approach proposed in Ref. [37], which identified 14 experimental spins using high-resolution data (around 10,000 time points of the coherence signal) across multiple 13C Larmor periods, our method identifies more spins using a much sparser dataset. The tradeoff of working with sparse data is the persistence of ill-posedness; however, our Bayesian inference framework reduces the number of plausible lattice sites from 3,518 to 230 by leveraging posterior detection probabilities.
@@ 6 Discussion
Crucially, the hybrid MCMC framework does not depend on the availability of DFT-derived hyperfine tensors and remains fully operational even in their absence. When the crystal lattice is known but detailed electronic-structure information is unavailable, the same hybrid sampling strategy can be applied over discrete lattice sites using magnetic dipolar hyperfine interactions computed analytically and directly from the fixed lattice geometry. This lattice-constrained formulation captures the dominant distance- and orientation-dependent physics governing the nuclear–electron interaction while substantially reducing the dimensionality of the inference problem. A continuous RWMH scheme can then relax the hyperfine coupling parameters associated with each lattice site around their approximate dipolar values, without relaxing the underlying lattice positions themselves. This hybrid inference strategy establishes the approach as a broadly applicable and practical technique for reconstructing nuclear spin environments across a wide class of spin-defect systems, including semiconductors for which reliable DFT hyperfine calculations are unavailable or impractical to compute. More generally, in systems where even lattice-resolved prior information is unavailable, the same Bayesian framework can be implemented in a fully continuous trans-dimensional setting using a continuous formulation of RJMCMC [27] or continuous RWMH sampling over both the number and spatial coordinates of nuclear spins without requiring any discrete lattice constraints.
@@
A potential limitation of the hybrid MCMC approach arises when the true Hamiltonian does not lie within the predefined class of candidate models. However, the framework is naturally compatible with established techniques for model selection and averaging, which can mitigate issues related to model misspecification. Bayesian model averaging [48], mixture models [49], and ensemble strategies [50] such as bagging offer systematic approaches to incorporate uncertainty over model structure. In particular, bagging methods—originally developed in the context of machine learning—construct ensembles by training models on bootstrapped subsets of the data and combining their predictions through voting, averaging, or regularized aggregation to improve robustness; such procedures are adopted in sparse data regimes where small variations in the dataset may lead to significant shifts in the inferred posterior [51]. By enabling principled treatment of both parameter and model uncertainty, such extensions could preserve the interpretability and physical grounding of the Bayesian approach while enhancing its stability and generalizability. Nonetheless, if the true model lies far outside the chosen candidate set, the current framework cannot generate a fully data-driven model; some choice of forward model or candidate class is always required. This feature distinguishes our approach from purely data-driven or machine learning–based methods. At the same time, leveraging substantial prior information about the physical system, such as Hamiltonian structure and DFT-based priors, is precisely what allows us to extract meaningful information from this highly ill-posed inverse problem.
Another potential limitation of MCMC-based strategies is their computational cost, which can become significant for high-dimensional models. This is an inherent feature of sampling-based approaches: the walkers require a burn-in period to explore the posterior distribution before producing reliable samples, and each likelihood evaluation can be expensive, particularly when forward models involve complex physics or first-principles calculations. While surrogate models [52] and reduced-order approximations [53] can mitigate the computational burden, the fundamental requirement for a sufficiently long burn-in and adequate sampling cannot be eliminated. High-dimensional problems require more computational resources, and careful attention must be paid to balancing model complexity, sampling efficiency, and available computational power.
@@ Acknowledgments
We thank F. Joseph Heremans, Mykyta Onizhuk, Daniel Sanz-Alonso, Nathan Waniorek, Christopher Egerstrom, Benjamin Pingault, and Christina Ioannou for useful discussions. A.N.P. acknowledges support from the DOE CSGF under Award Number DE-SC0022158. This research used resources of the University of Chicago Research Computing Center. This work was supported by AFOSR Grant No. FA9550-22-1-0370.
@@ Appendix A Hyperparameter dependence
Figure 6: Error trajectories of 5 ensembles as a function of walker steps for varying $\sigma^{2}$. The recovery algorithm was applied to the same coherence signal generated using $N=16$ CP pulses, $\varepsilon=0.001$ shot noise, hyperfine perturbation $0.001$, $\tau_{\text{max}}=0.008$ ms, and 250 interpulse times sampled from 15 simulated spins with $R_{\text{spin}}=5$ Å.
@@
We find that the hybrid algorithm presented is highly robust to $R_{\text{spin}}$, but the performance of the hybrid algorithm is sensitive to $\sigma^{2}$. In Fig. 7, we find that the variance, mean, and convergence properties of the error trajectories are uncorrelated with the size of the walker radius over the lattice sites $R_{\text{spin}}$. That is, the probabilistic selection rules introduced by the hybrid algorithm bias the walkers toward the same solution, regardless of the size of the neighborhood by which each lattice walker is constrained. Notably, there is a physical limit on how small the walker radius can be, which is the distance of nearest neighbors between lattice sites. For diamond, the distance between nearest neighbors is around $1.54$ Å. Unless otherwise stated, all simulations and experimental data recovery were performed with $R_{\text{spin}}=5$ Å.
@@
We note that hyperparameters need not be strictly optimal: beyond a certain point, further optimization yields diminishing returns, and the cost of hyperparameter optimization should not exceed the cost of the inference itself. Our primary requirement for hyperparameter values is reliable convergence of the MCMC walkers to the posterior distribution on an accessible timescale given available computational resources. In practice, tuning the hybrid algorithm involves balancing exploration and convergence across both discrete and continuous parameter spaces, although one should also account for the trade-off between optimizing hyperparameters, which can lead to faster convergence, versus the time-consuming and computationally intensive process of optimizing the hyperparameters.
For continuous parameters, adjusting proposal step sizes or covariance matrices can help achieve an acceptance rate that promotes efficient mixing while avoiding local trapping. For discrete parameters, such as lattice site selection, the choice of neighborhood radius should respect physical constraints (e.g., nearest-neighbor distances) but can otherwise be broad without degrading convergence. Parallel tempering parameters, including the number of strands and inverse temperatures, can be selected to ensure sufficient exchange between chains, particularly in multimodal or rugged likelihood landscapes. Finally, the noise variance $\sigma^{2}$ in the likelihood should be tuned to balance model flexibility against overconfidence: too small values risk trapping the walkers in local minima, whereas too large values reduce the algorithm’s ability to accurately recover underlying parameters. In many cases, adaptive strategies provide a practical approach to systematically improve performance without requiring exhaustive hyperparameter searches [54]; such as monitoring acceptance rates and effective sample sizes during burn-in and adjusting proposal distributions or tempering schedules accordingly.
@@ Appendix B Performance of hybrid algorithm in dense, low noise limit
Despite these favorable trends, performance in the 0–25 kHz regime remains substantially degraded in terms of accuracy of the number of spins detected (3̃0%) and false positive rate (4̃0%) as seen in Figure 8, even in the limit of long dynamical decoupling sequences, large numbers of dynamical decoupling pulses, and dense samplings of interpulse spacings. This behavior should not be interpreted as an algorithmic failure, but rather as a manifestation of the fundamentally ill-posed nature of the inverse problem associated with weakly coupled nuclear spins. In this regime, individual spins induce only minimal modulations of the NV coherence signal, and the resulting signal variations are often comparable in magnitude to experimental noise [55]. Thus, many distinct spin configurations, differing in both number and hyperfine couplings, produce coherence signals that are effectively indistinguishable within the accessible measurement resolution.
The ill-posedness is further exacerbated by the large number of nuclear spins at small hyperfine couplings, which leads to a combinatorial proliferation of nearly degenerate solutions in parameter space. Even with extensive data, including large numbers of sampled time points and high pulse numbers, the available measurements do not uniquely constrain the underlying spin configuration. In this setting, additional data primarily reinforce existing ambiguities rather than resolve them, reflecting a saturation of the information content of the coherence signal under fixed experimental parameters.
Within our Bayesian framework, this intrinsic non-identifiability is made explicit through the structure of the posterior distribution, which exhibits broad, multi-modal support in the weak-coupling regime. The strict detection criteria employed here, requiring consistency between detected spins and lattice-resolved hyperfine couplings, favor conservative reconstructions at the expense of reduced apparent detection rates for extremely weakly coupled spins. Importantly, this trade-off reflects a principled response to an underdetermined inverse problem rather than a limitation of the inference procedure itself.
These observations indicate that improving performance in the 0–25 kHz regime cannot be achieved solely through algorithmic refinement or increased data density under otherwise fixed experimental conditions. Instead, resolving such weakly coupled spins requires additional, complementary experimental information that alters the structure of the inverse problem itself, for example by combining measurements performed under different magnetic fields. Hence, the reduced performance in the weak-coupling regime reflects a fundamental information limit of the measurement setting, rather than a deficiency of the Bayesian inference framework.